OpenStudy (anonymous):

Solve 2x2 = 14x + 20

hartnn (hartnn):

don't know how to solve a quadratic equation ?

hartnn (hartnn):

can u factorize ? 2x^2-14x-20 = 0

OpenStudy (anonymous):

Yes I know how to factorize it, but the options given are confusing

OpenStudy (anonymous):

2x2 + 14x + 20 = 0 Reorder the terms: 20 + 14x + 2x2 = 0 Solving 20 + 14x + 2x2 = 0 Solving for variable 'x'. Factor out the Greatest Common Factor (GCF), '2'. 2(10 + 7x + x2) = 0 Factor a trinomial. 2((5 + x)(2 + x)) = 0 Ignore the factor 2. Subproblem 1Set the factor '(5 + x)' equal to zero and attempt to solve: Simplifying 5 + x = 0 Solving 5 + x = 0 Move all terms containing x to the left, all other terms to the right. Add '-5' to each side of the equation. 5 + -5 + x = 0 + -5 Combine like terms: 5 + -5 = 0 0 + x = 0 + -5 x = 0 + -5 Combine like terms: 0 + -5 = -5 x = -5 Simplifying x = -5 Subproblem 2Set the factor '(2 + x)' equal to zero and attempt to solve: Simplifying 2 + x = 0 Solving 2 + x = 0 Move all terms containing x to the left, all other terms to the right. Add '-2' to each side of the equation. 2 + -2 + x = 0 + -2 Combine like terms: 2 + -2 = 0 0 + x = 0 + -2 x = 0 + -2 Combine like terms: 0 + -2 = -2 x = -2

OpenStudy (anonymous):

I found my answer, I just had some trouble

hartnn (hartnn):

reordering is incorrect

hartnn (hartnn):

did u get how 2x^2-14x-20 = 0 ?

hartnn (hartnn):

2x2 = 14x + 20 subtracting 14x and 20 on both sides 2x^2-14x-20 =0 2(x^2-7x-10) =0 u need to solve x^2-7x-10 =0 using the quadratic formula.

hartnn (hartnn):

factoring will not help.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I had to find both x's

OpenStudy (anonymous):

-5 and -2

hartnn (hartnn):

-5 and -2 is incorrect because -5*-2 = 10 and not -10 factoring will not help here, u need to use quadratic formula